# Note on Using Singular Value Decomposition to Solve for Single-Phase Flow and Identify Isolated Clusters in Heterogeneous Pore Networks

### v3

@@ 3.2 Diluted pore networks
As shown in Figure 3a, the numerical values obtained for $p_{c}$ are in very good agreement with the theoretical values both for the hexagonal and square lattices. Slight deviation of the numerical from the theoretical values can be attributed to the finite size of the lattices with $N=100$, and the finite number of realizations with $N_{R}=50$. Figure 3b shows the dependence of permeability on $p_{b}$. For $p_{b}<p_{c}$, permeability tends to zero, then for $p_{b}>p_{c}$ it increases monotonically toward $k_{r}=1$, which is the value for the non-diluted lattices. For $p_{b}$ far from $p_{c}$, the evolution is approximately linear. These behaviors are in accordance with the literature [20, 21]. The variability of the $k_{r}$ values between realizations, illustrated by the width of the error bars, is not linear, with zero variance for $p_{b}=1$ and $p_{b}\ll p_{c}$ and a large variance for intermediate values. These results demonstrate the robustness of the proposed SVD method.
@@ 4 Conclusion
We have shown that SVD is a robust, easy-to-implement, and accurate method to identify sites belonging to isolated clusters in heterogeneous networks. Isolated clusters render the modified graph Laplacian non-invertible. Thus, we use the pseudoinverse determined from SVD to solve for single-phase flow in the heterogeneous networks, which assigns zero pressure to sites belonging to isolated clusters. The SVD method is demonstrated and validated for pore network models of different complexity, which are characterized by singular points and isolated clusters. This shows that SVD provides an alternative to iterative search algorithms for the identification of isolated clusters and a powerful tool to solve for single-phase flow on heterogeneous networks without pre-conditioning.

### v4

@@ 3.2 Diluted pore networks
From the SVD-based numerical flow simulations in the diluted networks, the percolation threshold $p_{c}$ is estimated as the bond occupation probability $p_{b}$ for which $50\%$ of the lattice realizations percolate [11]. As shown in Figure 3a, the numerical values obtained for $p_{c}$ are in very good agreement with the theoretical values both for the hexagonal ($p_{c}\approx 0.6527$) and square ($p_{c}=0.5$) lattices. A slight deviation of the numerical from the theoretical values can be attributed to the finite size of the lattices with $N=100$, and the finite number of realizations with $N_{R}=50$. Figure 3b shows the dependence of permeability on $p_{b}$. For $p_{b}<p_{c}$, permeability tends to zero, then for $p_{b}>p_{c}$ it increases monotonically toward $k_{r}=1$, which is the value for the non-diluted lattices. For $p_{b}$ far from $p_{c}$, the evolution is approximately linear. These behaviors are in accordance with the literature [17, 12]. The variability of the $k_{r}$ values between realizations, illustrated by the width of the error bars, is not linear, with zero variance for $p_{b}=1$ and $p_{b}\ll p_{c}$ and a large variance for intermediate values. These results demonstrate the robustness of the proposed SVD method.
@@ 4 Conclusion
We have shown that SVD is a robust, easy-to-implement, and accurate method to identify sites belonging to isolated clusters in heterogeneous networks. Isolated clusters render the modified graph Laplacian non-invertible. Thus, we use the pseudoinverse determined from SVD to solve for single-phase flow in the heterogeneous networks, which assigns zero pressure to sites belonging to isolated clusters. The SVD method is demonstrated and validated for pore network models, which are characterized by singular points and isolated clusters. The time complexity for algorithms to identify isolated clusters such as the one proposed in [22] is of order $N\ln N$, where $N$ is the number of sites, while the time complexity for matrix inversion is of the order of $N^{3}$. The proposed SVD method, which provides at the same time the flow solution and delineates the isolated clusters, is of order $N^{3}$. This shows that SVD provides a robust alternative to combined algorithms that identify isolated clusters and solve for single phase flow sequentially.